OpenStudy (anonymous):

Solve ln(3x-3)=1. round to the nearest thousand?

OpenStudy (mathstudent55):

\(\large \log_e(3x - 3) = 1\) Turn the log equation into an exponential equation by doing this: \(\log_b x = y \iff b^y = x\) \(e^1 = 3x - 3\)

OpenStudy (mathstudent55):

Remember that e^1 is simply e which is a number.

OpenStudy (anonymous):

@mathstudent55 im stuck what would i do next?

OpenStudy (shamil98):

\[\ 3x - 3 = e \] just solve for x.

OpenStudy (anonymous):

@shamil98 would it just be 1/3?

OpenStudy (shamil98):

no..

OpenStudy (shamil98):

solve the equation for x, then punch into a calculator since they want the exact form.

OpenStudy (anonymous):

@shamil98 1+1/3e ? lol thats what i got -_-

OpenStudy (shamil98):

\[3x - 3 = e\] \[\ 3x = e+3 \] \[\ x = \frac{e+3}{3} \] How did you get 1+1/3e ......

OpenStudy (anonymous):

@shamil98 i calculated wrong my bad! thank you :) cause i forgot what to do lol

OpenStudy (shamil98):

mhm.